jp:

What is the solution for the equation 4x − 10 = 2x? x

jp:

ok

Shadow:

\[4x - 10 = 2x\] Subtract 4x from both sides \[-10 = -2x\] Multiply both sides by -1 \[10 = 2x\] Divide both sides by 2 \[x = 5\]

glizxzy:

x = 1.666666667

Shadow:

Where'd you get that # from?

glizxzy:

alot of problem solivng

glizxzy:

would you like to see the equation

glizxzy:

Simplifying 4x = 10 + -2x Solving 4x = 10 + -2x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '2x' to each side of the equation. 4x + 2x = 10 + -2x + 2x Combine like terms: 4x + 2x = 6x 6x = 10 + -2x + 2x Combine like terms: -2x + 2x = 0 6x = 10 + 0 6x = 10 Divide each side by '6'. x = 1.666666667

Buenos88:

ummmmm that is not how you solve it and you ended up confusing me on your equation. and one side would be positive the other negative

jp:

lol

Buenos88:

yeah

jp:

oop

glizxzy:

no yall arent as smart XD

Buenos88:

you sure about that?

glizxzy:

iq

jp:

peeps

Shadow:

2x isn't negative

jp:

Buenos88 already explained the solution to me

Shadow:

Why didn't he do it in the question?

Shadow:

We usually prefer that so if others have the same question, they can also benefit from the help.

jp:

she did help me and explained it

jp:

oof

Buenos88:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Shadow Why didn't he do it in the question? \(\color{#0cbb34}{\text{End of Quote}}\) its she thank you very much. I am not trans yet.

jp:

lol

Buenos88:

thx

jp:

np